Vang Corp.'s stock price at the end of last year was $26 and its earnings per share for the year were $2.30. What was its P/E ratio?

Select the correct answer.

a. 12.50
b. 11.90
c. 13.10
d. 13.70
e. 11.30
Answer #1

e. 11.30

P/E ratio = MPS/EPS
= 26.00/2.30
=         11.30
Where,
MPS = Market price per share = $    26.00
EPS = Earning per share = $       2.30